expand the epression ln(5x/2)
condense the expression log4 3+3log4 2

Answer by edjones(8007)

ln(5x/2)=ln(5x)-ln(2)
.
log4 3+3log4 2=log[4](3*2^3)=log[4](24)
.
Ed